OpenStudy (anonymous):

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura returned the movie and paid $5.00, how many nights did she keep the movie? 14 1.5 8 6

OpenStudy (anonymous):

@JakeV8 this one is different!

OpenStudy (anonymous):

Yeah, it's different, but it's the same idea. You can do it... I'll help :)

OpenStudy (anonymous):

your equation should be 2+.5x=5

OpenStudy (anonymous):

do you know why?/

OpenStudy (anonymous):

no @tercespot

OpenStudy (anonymous):

Hm. Maybe I'll try again :)

OpenStudy (anonymous):

@broweyebrit , you need a cost function, like before... cost = flat fee plus (per night cost) * (number of nights) c(x) = flat fee + (per night fee) * x c(x) = 2 + 0.50x Does that make sense so far?

OpenStudy (anonymous):

that makes it seem a little more complicated that it is. since you have a definite fee of 2 dollars you have to add that to the equation.

OpenStudy (anonymous):

then you put x to represent how many night he had the movie extra

OpenStudy (anonymous):

and since it was 50 cents per night you multiply them together

OpenStudy (anonymous):

then add them together to get your overall payment

OpenStudy (anonymous):

so subtract 2 from the overall payment to isolate x

OpenStudy (anonymous):

then divide by .5 to know how many night he had the movie extra

OpenStudy (anonymous):

ughh im soo confused!!

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

ok sorrry, so you dont understand the equation?

OpenStudy (anonymous):

no @tercespot

OpenStudy (anonymous):

ok, so what are you struggling with?